### v1

# Embedded Hyperspectral Band Selection with Adaptive Optimization for Image Semantic Segmentation

Abstract
Hyperspectral band selection plays a pivotal role in remote sensing and image analysis, aiming to identify the most informative spectral bands while minimizing computational overhead. In this paper, we introduce a pioneering approach for hyperspectral band selection that offers an embedded solution, making it well-suited for resource-constrained or real-time applications. Our proposed method, embedded Hyperspectral Band Selection (EHBS), excels in selecting the best bands without the need for prior processing, seamlessly integrating with the downstream task model. This is achieved through the adaptation of the Stochastic Gates (STG) algorithm, originally designed for feature selection, for hyperspectral band selection in the context of image semantic segmentation and the integration of a dynamic optimizer, DoG, which removes the need for the required tuning the learning rate.
To assess the performance of our method, we introduce a novel metric for evaluating band selection methods across different target numbers of selected bands quantified by the Area Under the Curve (AUC). We conduct experiments on two distinct semantic-segmentation hyperspectral benchmark datasets, demonstrating its superiority in terms of its resulting accuracy and its ease of use compared to many common and state-of-the-art methods. Furthermore, our contributions extend beyond the realm of hyperspectral band selection. The adaptability of our approach to other tasks, especially those involving grouped features, opens up promising avenues for broader applications within the realm of deep learning, such as feature selection for feature groups. The demonstrated success on the tested datasets and the potential for application to a variety of tasks underscore the value of our method as a substantial addition to the field of computer vision. It offers new possibilities for feature selection and optimization in complex data analysis scenarios, making it a significant asset for researchers and practitioners alike.
1 Introduction
Hyperspectral imaging (HSI) involves capturing the complete optical spectrum at each point within an image. While a standard color camera records light intensity in just three colors (Red, Green, and Blue), a hyperspectral camera captures the entire wavelength range (typically consisting of several hundred bands) of light reflected from each scene point. The transition from color to full hyperspectral imaging provides a substantial increase in information, holding considerable potential across various applications such as medical imaging, agriculture, aerial photography, and autonomous driving.
While HSI offers notable advantages, it is not without its challenges. One significant drawback lies in the substantial costs associated with the physical sensor hardware. Additionally, HSI incurs increased expenses related to storing, transferring, and analyzing the considerably larger image data generated by HSI. Furthermore, the rise of very large and deep networks for vision has further increased the complexity of models and, consequently, computing costs. Hence, there is a pressing need for algorithms and methods for band selection, i.e., identifying a subset of hyperspectral bands that retains essential information for downstream tasks. Though existing research has explored band selection, the various methods proposed typically involved an unsupervised pre-processing step that is independent of the downstream HSI task and thus the choice of band selection may be sub-optimal.
This paper introduces Embedded Hyperspectral Band Selection (EHBS), a plug-and-play embedded method. EHBS effectively selects the optimal bands without requiring preliminary processing, seamlessly integrating with the downstream task model. EHBS utilizes an existing feature selection algorithm based on stochastic gates that was adopted to the setting of band selection of hyperspectral data. In this study, we showcase the capability of EHBS to dynamically learn optimal bands seamlessly as an integral part of the Convolutional Neural Network (CNN) implementation. This is done in the context of semantic segmentation, a visual task that predicts semantic categories for each pixel in an input image, that has received growing interest in the context of hyperspectral imagery, particularly with the application of deep learning methods. Our results demonstrate that our method outperforms existing band-selection methods on two different hyperspectral semantic segmentation datasets achieving similar accuracy values of the full hyperspectral data by using only around 25% of the bands. Importantly, our approach stands out by being easily applicable to deep-learning tasks over HSI datasets.
2 Background and Related Work
2.1 Hyperspectral Imaging
Hyperspectral imaging captures rich spectral data per pixel, unlike standard imaging with only three spectral samples (Red, Green, Blue). Contemporary hyperspectral systems offer hundreds of spectral bands spanning both visible and invisible spectra, effectively creating a three-dimensional cube composed of two-dimensional grayscale images. This detailed HSI data provides insights beyond the capabilities of regular RGB imaging that can be exploited by deep learning models.
Applications of HSI data are broad and range from medical tissue classifying where ill tissues can be found with non-invasive methods [18, 7] to non-destructive quality assessment of agricultural products [26], autonomous driving [5] and face recognition [31].
Semantic segmentation is a fundamental HSI task to predict the semantic categories for each pixel of an hyperspectral input image [25]. Semantic segmentation tasks based on aerial and satellite images play an important role in a wide range of applications [20]. In recent years, the successful application of deep learning (DL) in the field of computer vision (CV) has led to a surge of work applying DL methods for data semantic segmentation, resulting in notable achievements and significant advancements.
HSI algorithms set themselves apart from typical image processing methods due to variations in dataset size and data samples. Despite each sample carrying significantly more information, hyperspectral datasets are orders of magnitude smaller compared to RGB datasets. This limitation presents challenges, requiring models to be concise enough to manage high-dimensional data effectively without an abundance of training data.
Due to the aforementioned reasons, early research in hyperspectral imaging (HSI) concentrated on the application of traditional machine learning (ML) models, with a particular emphasis on Support Vector Machines (SVMs) [28]. As deep learning models demonstrated increasing success in addressing computer vision tasks over RGB data, there has been a notable surge in work to apply deep learning to HSI, with Convolutional Neural Networks (CNNs) in particular gaining prominence [20].
Deep learning applications for HSI became popular and prevailed in recent years. For example, Zhang et. al [37] used a 3D CNN with transfer learning for aero image segmentation while [29] explored unsupervised hyperspectral unmixing using autoencoders to classify hyperspectral images into 3 labels (Tree,Water,Rock).
In the field of agriculture, [16] apply a deep learning CNN model to estimate strawberry ripeness from hyperspectral images. Acknowledging the computational burden associated with processing the entire spectrum data, the authors employed a sequential feature selector to enhance computational efficiency by reducing the number of bands. Another application with image-level context is face recognition using hyperspectral data. In [31], A 2D CNN model classified the face class using a single band image selected by a majority voting algorithm. However, such an approach does not scale to tasks in which multiple bands are required as in material detection where spectral data is very important [29].
In conclusion, deep learning has emerged as a pivotal and widely adopted technique in hyperspectral imaging (HSI) applications. The abundance of spectral information captured by HSI necessitates the development of effective band selection methods, a crucial consideration for reducing the input size fed into deep learning models. This becomes particularly essential given the challenges posed by the high dimensionality of hyperspectral data, emphasizing the ongoing need for innovative approaches to handle the intricacies of this unique imaging modality.
2.2 Feature Selection
Feature selection (FS) methods identify the essential features needed by a machine learning system to perform a downstream task and can be categorized broadly as filter, wrapper, and embedded methods. Filter methods eliminate irrelevant features before model learning, using statistical relevance scores [6, 10]. Wrapper methods determine feature relevance based on model performance [23, 4], but their drawback lies in computational expense [30]. Embedded methods address this by simultaneously learning the model and selecting relevant features within a self-contained and single-train organism. For example, the widely recognized Least Absolute Shrinkage and Selection Operator (LASSO) [34] is an embeded FS algorithm that minimizes loss with an $l_{1}$ constraint but is limited to linear functions. Attempts to extend LASSO using neural networks face challenges with suboptimal gradient descent on $l_{1}$ regularized objectives [15].
STG [36] propose an embedded feature selection neural network scheme. The STG procedure is based on probabilistic relaxation of the $\ell_{0}$ norm of features, or the count of the number of selected features. The $\ell_{0}$-based regularization relies on a continuous relaxation of the Bernoulli distribution; such relaxation allows the STG model to learn the parameters of the approximate Bernoulli distributions via gradient descent. The STG framework simultaneously learns either a nonlinear regression or classification function while selecting a small subset of features and provides an information-theoretic justification for incorporating Bernoulli distribution into feature selection. Figure 1 illustrates the STG model in which the stochastic gates are attached to the $x_{i}$ input features, where the trainable parameter $\mu_{i}$ and a noise component $\epsilon_{i}$ control the choice of the gate being active or not ($z_{i}$).
Figure 1: Illustration of the STG model
2.3 Hyperspectral Band Selection
Band selection differs from traditional feature selection in that, while feature selection typically involves deciding whether to include or exclude individual features, band selection operates at a broader level, focusing on the inclusion or exclusion of entire bands. This distinction is particularly relevant in computer vision models with patches or convolution layers, where band selection essentially translates to feature selection for entire groups.
Band selection (BS) methods are categorized into two main groups. Unsupervised methods operate without using any annotations, relying on information criteria for internal information and dissimilarity from other bands. While unsupervised methods are suitable for unlabeled datasets, their lack of consideration for specific hyperspectral imaging (HSI) task performance may result in sub-optimal band choices. On the other hand, Supervised methods utilize annotated data in the selection process, tailoring band selection to the specific task and considering prediction model performance. This category further divides into before-train (finding the best bands before training the downstream model) and embedded in train (training the band selector as part of the downstream model, as in our method). Following we will describe six different methods that can be classified into five different families: Search, Ranking, Clustering, Sparsity and DL-based. These methods were implemented and used as baseline models to be compared to our proposed model.
Search methods in band selection explore the band space using tailored information criteria to identify bands suitable for a specific task. As an exhaustive search of all combinations is impractical, search-based methods apply efficient searching heuristics based on predefined criteria. One such method, LP [14], employs dissimilarity criteria and an incremental search approach. At each iteration, the least similar band is added to the previously selected bands, estimating candidates as linear combinations. The search criteria include similarity metrics such as Bhattacharyya distance, Jeffries–Matusita (JM) distance [21], or spectral angle mapping (SAM) [19].
Ranking-based methods prioritize bands based on variance, dissimilarity, or other metrics to select the most important bands. These approaches, categorized as unsupervised and supervised, assess bands’ distinctiveness and contribution to specific tasks. In unsupervised ranking, high information criteria, like those in MVPCA [11] and CBS-CEM [12], focus on variance and dissimilarity, respectively. The latter aims to minimize band correlation or dependence. In supervised ranking, methods like MMCA [11] and Mutual Information (MI) [17] leverage labeled data to construct task-oriented criteria, minimizing misclassification error and prioritizing bands associated with ground truth labels. These ranking-based band selection algorithms efficiently identify relevant bands for specific tasks.
Clustering-based methods in band selection involve grouping bands and selecting one representative from each group to minimize redundancy. Often used in conjunction with ranking, these methods aim to choose the highest-ranked representatives from clustered bands [35]. WaLuDI and WaLuMI [27] employ a distance metric based on information theory measures to assess similarity between bands. Their approach minimizes in-cluster variance while maximizing between-cluster variance. The dissimilarity measure calculates the distance between bands using probability functions based on gray-scale pixel values. For instance, the KL divergence is employed to quantify differences between probability distributions, enhancing the effectiveness of band clustering and selection strategies.
Sparsity-based methods employ sparsity constraints to represent each band slice image as a linear combination of other bands, aiming to identify the most influential combinations. The Iterative Sparse Spectral Clustering (ISSC) algorithm [33] introduces an innovative approach by representing band connections as a graph and identifying representatives from each cluster. ISSC minimizes non-zero elements in a similarity matrix, where each row contains coefficients of corresponding bands. The algorithm utilizes spectral clustering to create clusters and selects the closest band to each cluster center as a representative.
Another more recent line of work include DL-based methods aimed to model the nonlinear interdependencies between the various spectral bands. BS-Nets [9] is an embedded unsupervised method that applies a DNN to reconstruct the full HSI image from partially available bands. The networks attention mechanism was used from both spatial and spectral views to infer the best bands combination for reconstruction.
3 Method
3.1 Problem statement
Let $X$ represent a sample of m data instances where each instance is an n-sized array of 2D images and $Y$ denotes the corresponding m labels.
Let $F$ be a family of models for the downstream task each accompanied by a choice of parameters $\theta$ and $Loss$ is a loss function between a specific label $x_{i}$ and a corresponding model output $y_{i}$.
We denote a possible band selection via an indicator vector $I\in\{0,1\}^{n}$ where $I_{j}=1$ iff band j was chosen for processing. The norm $\lVert I\rVert_{1}$ of an indicator function $I$ corresponds to the number of selected bands. We denote $x\odot I$ as the point-wise product between an input item $x$ and the indicator vector $I$ in which all non selected bands are effectively masked to zero. Let $k$ be the target number of bands.
The goal of embedded band selection methods is to simultaneously select and Indicator vector $I$ and a model $f_{\theta}\in F$ that minimize the overall loss of the data as follows:
$\arg\min_{\theta,I}\frac{1}{m}\sum\limits_{i=1}^{n}(Loss(f_{\theta}(x_{i}\odot I),y_{i}))$ (1)
3.2 Our proposed system - EHBS
The Embedded Hyperspectral Band Selection (EHBS) system, our proposed approach, is an end-to-end embedded system. It comprises a downstream task model enhanced with an additional layer inserted between the input layer and the task model. This added layer, is based on the principles of the Stochastic Gates (STG) model - see 2.2. Our novel adaptation is specifically tailored for hyperspectral band selection within the context of image semantic segmentation and is further detailed in 3.2.1. By adding this layer, EHBS leverages the intrinsic characteristics of hyperspectral data and addresses the unique requirements of semantic segmentation tasks, seamlessly integrating with downstream models without the need for band-selection prepossessing.
3.2.1 Adapting STG for Feature Groups and Convolutional Layers
Figure 2: Illustration of our EHBS layer
In the setting of band selection, and in contrast to the standard feature selection setting, all features of a given band should either be included or excluded from the input. We have thus adapted the stochastic gates framework, originally designed for feature selection, to work over groups of features. This is done be altering the gates layer to either mask all of the features in the group (i.e. band) or to leave the features intact. This layer comes right after the input layer and precedes the first layer of the deep learning network of the downstream task. The Gate is applied to each band-specific 2D input in the corresponding full HSI 3D input. The output of the gate is a corresponding mask on the input based on each specific gate value.
The stochastic gates layer includes $N$ gates per each band ($N$ is the size of the spectral dimension). Each gate $s$ includes a learned parameter $\mu$ which corresponds to the significance of its corresponding input band (See Figure Figure 2). The gates layer acts as a sparse layer manipulating the input layer while preserving its shape. This is done by multiplying all corresponding feature of a band by a calculated value based on the gates $\mu$ value.
At every step of the deep learning forward pass, the gates layer masks the input band features based on the following gate-specific calculated value: $z=clamp(\mu+\epsilon)$ where $\epsilon$ is a normally distributed noise with mean 00 and standard deviation $\sigma$ ($\epsilon\sim\mathcal{N}(0,\sigma^{2})$) and $clamp$ constrains the value to be in the range $[0,1]$ ($clamp(x)=max(0,min(x,1))$).
The inclusion of noise enables model exploration during the training phase; however, it is omitted during the testing and production stages. As we want the gates mask to converge into 1 or 0 (effectively performing band selection), we add a regularization component $R$ to the downstream task loss function calculated as follows:
$R=\lambda\sum_{i=1}^{N}\Phi(\frac{\mu_{i}}{\sigma})$ (2)
Where $\Phi$ is the standard Gaussian CDF and $\mu_{i}$ is the $\mu$ value of gate $s_{i}$ and $\lambda$ is a regularization factor.
3.2.2 Integrating a parameter free optimizer
An additional significant aspect of EHBS is the integration of a dynamic optimizer named DoG, which eliminates the need for meticulous tuning of learning rates, a common requirement in embedded feature selection methods. This dynamic optimization strategy adeptly navigates the trade-off between feature selection step size and accuracy, contributing to a more refined and adaptive hyperspectral band selection process.
From our initial experimentation, it was evident that the STG-based deep learning network is very sensitive to the setting of the learning rate. Hence, one needs to experiment and set a different learning rate value depending on the system architecture, input data and target number of bands. In order to circumvent this limitation we chose DoG, a parameter-free stochastic optimizer DoG (”Distance over Gradients”) [22]. The DoG step sizes depend on simple empirical quantities (distance from the initial point and norms of gradients) and there is no “learning rate” parameter that needs to be set.
We have used this dynamic optimizer to coup with the need to balance between the band selection objective to the segmentation loss objective in the combined training process. Employing a parameter-free stochastic method such as DoG has relieved us from the necessity of tuning the learning rate parameter thus providing a consistent model applicable to all experimental settings.
4 Experiments setting
4.1 Benchmark Datasets for Evaluation
In this section, we present the benchmark datasets employed to evaluate the performance of our proposed EHBS model for Image Semantic Segmentation. We focus on two common hyperspectral semantic segmentation benchmark datasets: Pavia University (PaviaU) and Salinas that represent diverse real-world scenarios.
4.1.1 PaviaU
The PaviaU hyperspectral semantic segmentation benchmark dataset [2] captures an urban scene over Pavia, northern Italy, acquired by the ROSIS sensor during a dedicated flight campaign. This publically available dataset encompasses 103 spectral bands in wave length of 430-860 nm, and the image has dimensions of 610 by 610 pixels. It is noteworthy that certain samples in the image lack information, resulting in 42,000 valid pixels for analysis. The dataset is annotated with ground-truth labels, classifying nine distinct categories, including bitumen, asphalt, tiles, trees, and more. These annotations provide a rich foundation for evaluating the semantic segmentation performance of models across a diverse set of urban materials and features. Figure 3 displays a sample band of the scene in grey-scale and as as the corresponding color-coded annotation of the ground-truth.
Figure 3: Pavia University dataset: (a) A sample band of scene in grey-scale and (b) The color-coded annotation of the ground-truth
4.1.2 Salinas
The Salinas hyperspectral semantic segmentation benchmark dataset [3] captures an agricultural scene in the Salinas Valley, California, gathered by the 224-band AVIRIS sensor in wave length of 430-2500 nm. Renowned for its high spatial resolution (3.7-meter pixels), the dataset covers an extensive area spanning 512 lines by 217 samples, providing approximately 54,000 valid pixels for analysis. The dataset is meticulously annotated to distinguish among 16 classes of ground truths, including vegetables, bare soils, and vineyard fields. Notably, to refine the dataset for agricultural semantic segmentation tasks, 20 water absorption bands were selectively discarded from the original dataset, specifically bands [108-112], [154-167], and 224.
Figure 4: Salinas dataset: (a) A sample band of the scene in grey-scale and (b) The corresponding color-coded annotation of the ground-truth
Figure 4 displays a sample band of the scene in grey-scale and as as the corresponding color-coded annotation of the ground-truth.
4.2 Band Selection Benchmark Methods
We selected seven different baseline methods for comparison with our proposed EHBS method in the context of embedded hyperspectral band selection for image semantic segmentation. Five methods, namely LP [14], ISSC [33], WALUMI [27], WALUDI [27], and MMCA [11], were identified as top-performing techniques in a comparative study [32] conducted on the Pavia dataset. We also included BS-Nets [9], an embedded unsupervised DL-based BS method. These supervised and unsupervised methods collectively represent five distinct band selection family types (See 2.3), offering a diverse and comprehensive benchmark for evaluating the performance of our proposed techniques.
Although not used before for band selection, we have implemented a supervised deep learning embedded method, where $\ell_{1}$ regularization was used on bands gates values, and the k gates with the highest values were the selected ones. This implementation was inspired by similar work in which regularization was embedded in DL networks for feature selection (see 2.2) and adapted to the setting of BS by applying the regularization over the bands rather than over the individual input features. The subsequent analysis and comparison against these established methods aim to provide a robust assessment of the efficacy and innovation introduced by our proposed embedded hyperspectral band selection methods.
4.3 Model Details
Our EHBS method is easily adaptable for integration into any CNN [24], ViT [13] or other deep learning model. In practice it is a neural layer between the input and the downstream task model, which consists of gates multiplying the input values with a learned factor for every patch input data proceeded by an arbitrary downstream task model. In our comprehensive experiments, we specifically focused on the state-of-the-art CNN model proposed by Hamida et el. [8], a 3D CNN specifically tailored for spectral-spatial data extraction. We chose this model for its proven performance in hyperspectral semantic segmentation tasks. The objective of our experiments was to evaluate the effectiveness of our embedded band selection approach in comparison to non-embedded methods across varied dataset sizes, sample (patch) sizes, and cross-validation scenarios. Implementation wise, we based our code on the publicly available implementation of Hamida et el. [8] available in [1]. The corresponding architecture of the CNN network is illustrated in Figure 5. The CNN network consists of several 3D convolutional layers, proceeded by a 1D convolution and eventually followed by a final class probabilities output layer. For the hyperparameters and kernel sizes, we used the default settings consistent with the original paper and its corresponding available model implementation  [8]. We used the Pytorch deep learning framework. Batch size was set to 256. The number of epochs was set to 100 for PaviaU and 150 for Salinas based on observing the converge patterns during our initial experimentation. The initial $mu$ and $sigma$ values of the stochastic gates layer were set to the recommended value of 0.5.
We utilized the DoG optimizer, eliminating the need for explicit learning rate tuning, contributing to the adaptability and efficiency of our method.
The STG layer includes a regularization component designed to minimize the number of the active gates (or more specifically their $\mu$ values). The corresponding regularization factor ($\lambda$) was used to meet the target number of bands. We run multiple experiments with various $\lambda$ values to obtain the various required target number of bands.
Figure 5: Illustration of Hamida et el. model
4.4 Experimental Design and Evaluation
In order to test our method and the benchmark methods in a robust and accurate way, we ran each experiment with 10-fold cross-validation. In addition, as most of the baseline methods are not embedded, we first run the various band selection models as a first stage and then used the selected bands from each method to train the downstream task model via the baseline semantic segmentation CNN model. This ensured a proper apples-to-apples comparison of the various band selection methods.
For evaluation, we compared the accuracy of the various methods for various target number of bands. In addition, in order to compare methods across the whole range of numbers of selected bands, we introduced an new evaluation metric which calculated the area under the bands-performance curve. This metric was inspired by the common Area Under the Curve (AUC) metric used to evaluate the performance of a binary classification model, particularly in the context of a Receiver Operating Characteristic (ROC) curve. The AUC provides a single scalar value that summarizes the performance of a classification model across various classification thresholds and is widely used metric for evaluating and comparing binary classification models. Similarly, in our setting, the AUC provides a single scalar value that summarizes the performance of a band selection model across various number of bands selected allowing us to compare different models by a single value.
For testing our EHBS model, in each experiment we run the model multiple times with different regularization factor ($\lambda$) values in the range 0.2 to 2.4 applying a simple heuristic search till the desired target number of selected bands was met. The number of bands that were used and the corresponding accuracy results were saved for each run. Overall accuracy for a given target number of bands was calculated by averaging the accuracy across all folds.
5 Results
Figure 6: Illustration of the selected bands for the Salinas dataset across various target number of bands. Each horizontal ”row” represents selection of bands in a single run of EHBS with the selected bands marked in black. The top row corresponds to the selection of 8 bands and the bottom row to a run with 48 bands selected.
(a) Accuracy results on the PaviaU dataset for the various methods
Figure 7: Accuracy and AUC results over the PaviaU dataset, 7x7 patch size, for multiple target number of bands
Bands Pavia 7x7 Pavia 11x11
Methods
6 (6%)
10 (10%)
15 (15%)
AUC
EHBS Waludi Walumi ISSC
BS-NETS
Conv
BS-NETS
FC
94.16 95.09 93.72 95.56 95.94 93.71
97.69 96.36 97.52 97.25 96.18 94.62
98.75 98.00 98.14 97.59 98.72 96.17
.969957 .965914 .966614 .968892 .969466 .949135
EHBS Waludi Walumi ISSC
BS-NETS
Conv
BS-NETS
FC
98.28 98.57 98.34 98.80 97.55 95.86
98.94 98.59 99.22 98.48 98.38 98.16
99.85 98.91 99.52 99.38 99.74 98.58
.990078 .986985 .990042 .988928 .986285 .976900
Table 1: Comparison Table - accuracy and AUC over the PaviaU dataset for two patch sizes (7x7, 11x11)
Bands Salinas 7x7 Salinas 11x11
Methods
10 (5%)
20 (10%)
30 (15%)
AUC
EHBS Waludi Walumi ISSC
BS-NETS
Conv
BS-NETS
FC
96.09 91.02 90.40 94.54 93.96 93.92
97.21 93.30 91.82 94.94 94.54 94.50
97.58 94.83 93.95 96.73 96.34 95.37
.970225 .931125 .919975 .952875 .94845 .945725
EHBS Waludi Walumi ISSC
BS-NETS
Conv
BS-NETS
FC
97.83 91.77 93.33 93.28 95.76 96.48
98.93 92.25 95.88 98.14 96.74 96.51
99.41 95.85 95.25 98.95 97.29 96.00
.98775 .9303 .95085 .97992 .966375 .96375
Table 2: Comparison Table - accuracy and AUC over the Salinas dataset for two patch sizes (7x7, 11x11)
Figure 7 shows a graph comparing the accuracy of our proposed model, EHBS, as well as the various baseline band selection models on the PaviaU dataset with a patch size of 7x7 over multiple target number of bands. As can be seen from the graph our proposed model outperforms the baseline models in most of the data points and is comparable to the best baseline model in the others. it also shows the corresponding AUC metric for the various models. Overall, our EHBS model achieved the highest AUC value surpassing the performance of all other models. Figure 8 shows the accuracy results of our proposed method over the whole range of target band selection for the PaviaU dataset and a 7x7 patch size. For this specific dataset, close to optimal performance is obtained by using the 20 best-selected bands (or $\sim$20of the bands). Performance is stable with only slight improvement from this point till the use of the full hyperspectral input.
Table 1 shows a comparison table including accuracy results and AUC results over the PaviaU dataset over two different patch sizes (7x7 and 11x11) and selected target number of bands. As baseline models we chose the top performing models on the PaviaU 7x7 experiments. Table 2 shows a similar comparison table for the Salinas dataset. For all datasets and patch sizes, EHBS obtained the highest AUC score and obtained the best accuracy for all Salinas experiments and for the vast majority of PaviaU experiments. In the few PaviaU experiments for which EHBS was not the top performing method it obtained comparable accuracy results to the top performing method.
Figure 6 illustrates the selected bands for the Salinas dataset across various target number of bands. EHBS systematically ignores areas with low contribution for the downstream task performance and the prominent band selection areas remain consistent across different runs.
Figure 8: EHBS accuracy result with Pavia 7x7 patch over the whole range of hyperspectral selection
6 Discussion and Conclusion
Following, we highlight several key findings and implications derived from our experiments. Firstly, our proposed method demonstrates clear superiority across two diverse datasets and various experimental settings, underscoring its robustness and effectiveness. We conducted comparisons among the methods in various settings, including different patch sizes, datasets, and the number of selected bands. We observed that for larger patch sizes, our method outperformed others by an even wider margin. We attribute this phenomenon to the presence of more contextual information and a higher model complexity. Consequently, employing an embedded model enables the model to discover the optimal bands that effectively capture the intricate non-linear interactions among the features. Additionally, our findings suggest that our model tends to outperform baseline models when confronted with an abundance of training data and a higher number of bands to select. We thus hypothesize that our method would perform even better on larger datasets, with larger patch sizes, and with different architectures that hold a global view per band. However, exploration of these scenarios will be reserved for future work. We demonstrated the strength of our model while integrated into CNN-based deep learning model for hyperspectral semantic segmentation. The seamless integration of our method into an existing deep learning model that was achieved by simply adding a selection layer after the input enhances its practicality and ease of use. This simplicity of implementation coupled with the robust performance positions our method as a promising and practical solution for various machine learning and computer vision tasks.
These results also demonstrate the robustness of our proposed method along multiple settings and various target number of bands. The method performs well and consistently without the need to tweak or modify the model regardless of the target number of bands - be it only a few number of bands or when aiming for a large number of selected bands. This also allows researchers to get a good understanding of the performance curve and choose the desired tradeoff given the application setting.
After our initial experimentation, we further tested more optimizers applying an extensive search for an optimal learning rate. As can be seen in Table 3, when using the Adam optimer the results are sensitive to the learning rate and finding the right learning rate is crucial. It turns out that on the PaviaU dataset with a 7x7 path, EHBS with the Adam optimizer can achieve even better performance than EHBS with DoG. However, this is a retrospective result obtained over our testset and not via tuning on an held out evaluation set. Given that the results achieved with the DoG optimizer are comparable to the optimal outcomes obtained using the Adam optimizer, we advocate for the adoption of DoG. Notably, DoG eliminates the need for parameter tuning, ensuring a more robust and reliable performance.
Bands DoG Adam lr=0.01 Adam lr=0.002 Adam lr=0.001
6 (6%) 94.16 76.67 98.97 98.20
10 (10%) 97.69 96.61 99.32 98.47
15(15%) 98.75 97.39 99.55 98.77
Table 3: Comparing accuracy of EHBS with different optimizers and learning rates (lr) for PaviaU dataset with patch size of 7
In conclusion, our embedded hyperspectral band selection method, leveraging the Stochastic Gates (STG) algorithm and the dynamic optimizer DoG, proves to be a promising solution for image semantic segmentation tasks. The seamless integration into downstream models, absence of pre-processing requirements, and adaptability to resource-constrained or real-time applications mark the method’s practical significance. Our approach, validated on two benchmark datasets, not only outperforms common and state-of-the-art methods in terms of information preservation but also introduces a novel AUC-based metric for evaluating band selection methods. The success of our method extends beyond hyperspectral band selection, showcasing its potential in broader applications within the domain of deep learning. The demonstrated efficiency and adaptability highlight its substantial contribution to computer vision, offering valuable possibilities for feature selection and optimization in diverse data analysis scenarios. Researchers and practitioners stand to benefit significantly from the versatility and performance of our proposed method.
7 Acknowledgement
This work was partially supported by the Chief Scientist of the Israeli Ministry of Agriculture grant number12-03-0010. We would also like to thank Nadav Spitzer for his help in implementing the baseline feature selection models and to Ofir Lindenbaum and Jacob Goldberger for providing input and feedback on our work.